# Growing Transformers: Modular Composition and Layer-wise Expansion on a Frozen Substrate

### v1

@@ Abstract
The prevailing paradigm for scaling large language models (LLMs) involves monolithic, end-to-end training, a resource-intensive process that lacks flexibility. This paper explores an alternative, constructive approach to model development, built upon the foundation of non-trainable, deterministic input embeddings. In prior work [1], we established that high-level semantic reasoning can emerge in Transformers using frozen embeddings derived from the visual structure of Unicode glyphs. Here, we demonstrate that this fixed representational substrate acts as a universal "docking port," enabling two powerful and efficient scaling paradigms: seamless modular composition and progressive layer-wise growth.
First, we show that specialist models trained on disparate datasets (e.g., Russian and Chinese text) can be merged into a single, more capable Mixture-of-Experts (MoE) model, post-training, with zero architectural modification. This is achieved by simply averaging their output logits. The resulting MoE model exhibits immediate performance improvements on reasoning benchmarks like MMLU, surpassing its constituent experts without catastrophic forgetting. Second, we introduce a layer-wise constructive training methodology, where a deep Transformer is "grown" by progressively stacking and training one layer at a time. This method demonstrates stable convergence and a clear correlation between model depth and the emergence of complex reasoning abilities, such as those required for SQuAD.
Our findings suggest a paradigm shift from monolithic optimization towards a more biological or constructive model of AI development, where complexity is built incrementally and modules can be composed freely. This opens new avenues for resource-efficient scaling, continual learning, and a more democratized ecosystem for building powerful AI systems. We release all code and models to facilitate further research.
INDEX TERMS constructive learning, continual learning, frozen embeddings, mixture of experts (MoE), model merging, modular AI, progressive learning, Transformers
@@ 1.   INTRODUCTION
In prior work [1], we challenged a core assumption of this paradigm: the necessity of trainable input embeddings. We demonstrated that Transformers can achieve robust convergence and strong reasoning performance using a completely frozen embedding layer based on deterministic, non-semantic visual representations of Unicode characters. This finding established that semantic understanding is an emergent property of the Transformer’s compositional architecture, not a feature of its input vectors.
This paper investigates the profound architectural implications of that discovery. If the input representation is a fixed, universal constant, can it serve as a standardized substrate for building models in a more modular and efficient manner? We answer this question affirmatively by proposing and validating a "constructive learning" framework. We hypothesize that a shared, frozen embedding space can act as a universal interface, or a "docking port," allowing separately trained neural components to be combined and complex architectures to be grown incrementally.
We validate this hypothesis through two sets of experiments:
  • Seamless Modular Composition: We demonstrate that specialized models, pre-trained on different languages, can be merged into a Mixture-of-Experts (MoE) model post-hoc. Because they share an identical representational input/output space, their learned knowledge can be combined by simply averaging their predictions, leading to a system that is more capable than its individual parts.
  • Progressive Layer-Wise Growth: We introduce a method for "growing" a Transformer model from the ground up. We start with a single Transformer layer and train it to convergence. We then freeze it, stack a new layer on top, and repeat the process. This incremental, layer-by-layer construction mimics a more natural growth process and allows us to observe the emergence of capabilities as a direct function of model depth.
Our results show that these methods are not only viable but highly effective, offering a path towards more sustainable, adaptable, and interpretable model scaling.
Code and all artifacts are released to foster further progress in both science and industry [^1] Repository: https://github.com/AVBochkov/Embeddings (https://huggingface.co/Bochkov).
@@ 2.1.   Greedy Layer-Wise Training
The idea of training deep networks one layer at a time was pioneered by Hinton et al. [2] for Deep Belief Nets and explored by Bengio et al. [3] for deep autoencoders. These methods were primarily used for pre-training to find a good initialization for subsequent end-to-end fine-tuning. Our approach differs fundamentally: we use progressive layer-wise training not for initialization, but as the primary, constructive training process itself, building upon a fixed, non-trainable base.
2.2.   Mixture of Experts (MoE) and Model Merging
MoE models, such as those by Shazeer et al. [4] and Fedus et al. [5], use gating networks to dynamically route inputs to specialized subnetworks. However, these are typically trained jointly from the start. Model merging techniques, such as Model Soups [6], often average model weights, which requires careful alignment. Our approach is radically simpler: because our expert models share an identical, deterministic input/output vocabulary mapping via the frozen embeddings, their logits are directly comparable. This allows for a "zero-cost" merge by simple averaging, a capability not present in standard LLMs with independently trained embeddings.
2.3.   Progressive and Modular Architectures
Progressive Neural Networks (PNNs) [7] achieve continual learning by adding new network "columns" for each new task. Adapter-based methods like AdapterFusion [8] inject small, trainable modules into a frozen base model. Our layer-wise growth is a form of vertical, rather than horizontal, expansion. The use of Low-Rank Adaptation (LoRA) [9] in our later growth stages serves as an efficient tool for global readjustment, but the core principle remains constructive and layer-wise.
@@ 3.   CONSTRUCTIVE LEARNING METHODOLOGY
The cornerstone of our methodology is the frozen visual Unicode embedding layer, as detailed in our previous work [1]. This provides a fixed, deterministic mapping from any token in a vocabulary to a $d_{\text{model}}$dimensional vector. This shared (V, $d_{\text{model}}$) embedding matrix is the immutable foundation for all models in this study.
3.1.   Seamless Model Composition via a Shared Substrate
Given two or more independently trained "expert" Transformer models, $M_{a}$, $M_{b}$, …, which were all trained using the identical frozen embedding layer, we can compose them into an MoE model $M_{\text{moe}}$. Each expert $M_{i}$ has its own trained Transformer blocks but shares the same input nn.Embedding layer and output projection shape.
For a given input sequence idx, each expert produces a logits tensor: $logits_{a}=M_{a}(\texttt{idx})$ and $logits_{b}=M_{b}(\texttt{idx})$. Since the vocabulary and its indexing are identical across models, the $j$-th element of the logit vector from any model corresponds to the same token. This enables a simple and effective merging strategy:
Logit Averaging: The combined logits are computed as $logits_{\text{moe}}=(logits_{a}+logits_{b})/2$. This method is parameter-free and can be applied at inference time without any further training.
Adapter-based Fusion: For potentially higher performance, the logits can be concatenated and passed through a small, trainable adapter: $logits_{\text{moe}}=\text{adapter}(\text{concat}(logits_{a},logits_{b}))$.
In our experiments, we primarily use logit averaging and demonstrate that even this simple approach yields significant gains, which can be further improved with minimal fine-tuning of the output layer.
@@ 3.2.   Progressive Layer-Wise Growth
Initialization: Create a model $M_{1}$ with a single Transformer block ($n_{\text{layer}}=1$, model: ’abs-bvv-1’) on top of the frozen embedding layer.
@@
Train Layer 2: Train $M_{2}$ (’abs-bvv-2’), where only the weights of the newly added second layer are trainable.
@@
For deeper models (e.g., $n_{\text{layer}}>=3$), we explore a variant where, upon adding a new layer, we use LoRA [9] to perform a low-rank update on the weights of all existing layers. This allows the entire network to adapt to the new depth with minimal trainable parameters, preventing ossification and promoting global coherence.
@@ 4.   EXPERIMENTS AND RESULTS
4.1.   Results: Seamless Model Composition (MoE)
We use the model families (best_bvv, max_bvv, nemo_bvv) from our previous work [1] for the MoE experiments. For the progressive growth experiment, we construct a new, larger model with $d_{\text{model}}=4096$ and $n_{\text{head}}=32$. All models use the frozen visual embeddings based on the bvv241 tokenizer. Training data is a $\approx$9B token mix of Wikipedia and SFT datasets. Evaluation is performed on MMLU, ARC, CommonsenseQA, and SQuAD.
We merged language-specific experts (EN + Russian and EN + Chinese) into MoE models. Critically, the training process for the MoE model did not start from a high-loss state. For instance, ‘best_bvv_moe‘ ’s initial validation loss was already low (2.7), close to the final loss of the experts, before converging further to 2.044 (Figure 1). This confirms that the composition immediately leverages the combined competence of the experts without catastrophic interference. Figure 2 shows a similar trend for ‘max_bvv_moe‘. Figure 3 visualizes the performance gains of the merged models over their constituent experts.
Figure 1: Training dynamics for the ‘best_bvv_moe‘ model. The low starting loss indicates successful knowledge transfer from expert models, followed by further optimization.
Figure 2: Training dynamics for the ‘max_bvv_moe‘ model, mirroring the successful convergence pattern seen in ‘best_bvv_moe‘.
Figure 3: Performance comparison: The merged MoE model (best_bvv_moe) demonstrates synergistic capabilities, achieving superior performance on key reasoning benchmarks like MMLU compared to its individual expert counterparts.
@@ 5.1.   The Frozen Embedding as a Universal Docking Port
The key enabler for our methods is the shared, frozen representational substrate. It acts as a universal standard, an "API" or "docking port" for neural components. When all models speak the same fundamental input/output language of visual forms, their higher-level learned knowledge (the transformations within their layers) becomes interoperable. This resolves a major hurdle in model merging and modularity, which is typically plagued by incompatible, independently learned embedding spaces.
@@ 5.3.   Implications for a Modular AI Ecosystem
  • Resource Efficiency and Specialization: Organizations could train smaller, expert models on proprietary or specialized data. These experts could then be merged or sold as compatible "plugins" for larger, general-purpose models.
@@ 6.   CONCLUSION
Building on the foundation of frozen visual embeddings, we have demonstrated two novel and efficient paradigms for scaling Transformer models: seamless post-hoc composition of expert models and progressive layer-wise growth. Our experiments show that specialist models can be merged into a more capable whole without costly retraining, and that deep models can be "grown" incrementally, with complex reasoning abilities emerging as a direct function of depth.
This work reframes the challenge of scaling AI from a monolithic endeavor to a constructive and modular one. By establishing a fixed, universal representational substrate, we unlock a new design space for creating powerful, flexible, and efficient AI systems. This "constructive learning" paradigm offers a promising path toward a more sustainable and collaborative future for artificial intelligence.

### v2

@@ Abstract
The prevailing paradigm for scaling large language models (LLMs) involves monolithic, end-to-end training, a resource-intensive process that lacks flexibility. This paper explores an alternative, constructive scaling paradigm, enabled by the principle of emergent semantics in Transformers with frozen, non-semantic input embeddings. We posit that because high-level meaning is a compositional property of a Transformer’s deep layers, not its input vectors, the embedding layer and trained lower layers can serve as a fixed foundation. This liberates backpropagation to focus solely on newly added components, making incremental growth viable. We operationalize this with a layer-wise constructive methodology that combines strict layer freezing in early stages with efficient, holistic fine-tuning of the entire model stack via low-rank adaptation (LoRA) as complexity increases. This method not only demonstrates stable convergence but also reveals a direct correlation between model depth and the emergence of complex reasoning abilities, such as those required for SQuAD, which are absent in shallower models. In a controlled study, our constructively grown model rivals the performance of a monolithically trained baseline of the same size, validating the efficiency and efficacy of the approach. Our findings suggest a path towards a paradigm shift from monolithic optimization towards a more biological or constructive model of AI development. This opens a path for more resource-efficient scaling, continual learning, and a more modular approach to building powerful AI systems. We release all code and models to facilitate further research.
@@ 1.  INTRODUCTION
A recent finding challenged a core assumption of this paradigm by demonstrating that Transformers can achieve robust reasoning performance using a completely frozen, non-semantic embedding layer (Bochkov, 2025). This established that semantic understanding is an emergent property of the Transformer’s compositional architecture, not an inherent feature of its input vectors.
This paper investigates a direct architectural consequence of this discovery: if semantics emerge from depth, can we build powerful models by progressively increasing that depth, layer by layer, upon a stable, frozen foundation? We answer this question affirmatively by proposing and validating a "constructive learning" framework for scaling Transformers. We hypothesize that a frozen embedding layer acts as an immutable substrate, allowing deeper computational structures to be built and trained incrementally, a process that is far more resource-efficient than monolithic training.
To validate this hypothesis, we conduct two distinct sets of experiments:
  • Feasibility at Scale under Resource Constraints: First, we demonstrate the methodology’s viability for building models whose full optimizer state exceeds the memory of a single accelerator. We progressively grow a large-scale Transformer to a final size of 2.3B parameters, while keeping the number of trainable parameters at each growth stage within a fixed budget (approx. 740M, fitting on a single H100 GPU). The training throughput was scaled using data parallelism across additional accelerators. This experiment shows how complex capabilities emerge as the model grows from 1 to 6 layers, demonstrating a practical path to training models that would otherwise be infeasible to train monolithically on such hardware.
  • Controlled Comparative Study: Second, to rigorously evaluate the performance of our approach, we conduct a controlled experiment on smaller (247.6M), 9-layer models. We compare our constructive method using frozen embeddings against three critical baselines: (a) a traditional, monolithically trained model of the same final architecture; (b) a constructively grown model using standard trainable embeddings; and (c) an ablation with a minimalist embedding composed of 16-component binary vectors to test the limits of the principle.
Our results show that this constructive approach is effective, offering a path towards more sustainable, adaptable, and interpretable model scaling. All code and models are released to foster further progress [^1] Repository: https://github.com/AVBochkov/PGT ( https://huggingface.co/Bochkov ).
@@ 2.1.  Greedy Layer-Wise Training
The idea of training deep networks one layer at a time was pioneered by Hinton et al. (Hinton et al., 2006) for Deep Belief Nets and explored by Bengio et al. (Bengio et al., 2007) for deep autoencoders. These methods were primarily used for pre-training to find a good initialization for subsequent end-to-end fine-tuning. Our approach differs fundamentally: we use progressive layer-wise training not for initialization, but as the primary, constructive training process itself, building upon a fixed, non-trainable base.
$2.$  Progressive and Modular Architectures
Progressive Neural Networks (PNNs) (Rusu et al., 2016) achieve continual learning by adding new network "columns" for each new task. Adapter-based methods like AdapterFusion (Pfeiffer et al., 2021) inject small, trainable modules into a frozen base model. Our layer-wise growth is a form of vertical, rather than horizontal, expansion. The use of Low-Rank Adaptation (LoRA) (Hu et al., 2022) in our growth stages serves as an efficient tool for global readjustment  (Schulman and Lab, 2025), but the core principle remains constructive and layer-wise.
@@ 3.  CONSTRUCTIVE LEARNING METHODOLOGY
The cornerstone of this methodology is the frozen visual Unicode embedding layer, as detailed in  (Bochkov, 2025). This provides a fixed, deterministic mapping from any token in a vocabulary to a $d_{\text{model}}$ -dimensional vector. This shared (V, $d_{\text{model}}$) embedding matrix is the immutable foundation for all models in this study.
@@ 3.1.  Progressive Layer-Wise Growth
Initialization: Create a model $M_{1}$ with a single Transformer block ($n_{\text{layer}}=1$) on top of the frozen embedding layer.
@@
Train Layer 2: Train $M_{2}$, where only the weights of the newly added second layer are trainable.
@@
This layer-wise process can be strategically interspersed with holistic fine-tuning stages. For deeper models (e.g., after reaching 3 or 5 layers), we introduce an alternative training phase. Instead of adding a new layer, we freeze the entire stack and apply Low-Rank Adaptation (LoRA) (Hu et al., 2022) adapters to all existing layers.
Crucially, this switch in strategy operates under a constant training budget. The number of trainable parameters in the LoRA adapters is carefully chosen to match the number of parameters in a single new Transformer layer ($\approx$740M in our large-scale experiment). This allows us to reallocate the same computational resources from growing the model’s depth to enhancing its internal coherence. This LoRA-based fine-tuning allows the entire network to adapt to its new depth, promoting global integration of knowledge and preventing the ossification of lower layers, all without increasing the instantaneous memory requirements for training.
4.  EXPERIMENTAL SETUP
4.1.  Overview of Experiments
To validate our constructive learning framework, we conduct two distinct sets of experiments. The first demonstrates the feasibility of our method for training large-scale models under typical hardware memory constraints. The second provides a rigorous controlled comparison of our approach against key baselines to evaluate its performance and isolate the contribution of frozen embeddings.
4.2.  Model Architectures and Baselines
All models in our experiments are standard decoder-only Transformers, similar to GPT-2, using rotary embeddings and GELU activations.
4.2.1.  Large-Scale Feasibility Study
For the feasibility study, we use a large model configuration with a hidden dimension $d_{\text{model}}=4096$, $n_{\text{head}}=32$, and a vocabulary size of 131,072. The model, which we refer to as ABS-BVV, is grown progressively from 1 to 6 layers. At each stage, only the newly added layer(s) are trained, keeping the number of trainable parameters constant at approximately 740M. This ensures the optimizer states fit within the memory of a single 80GB H100 GPU. The final 6-layer model reaches 2.3B parameters.
4.2.2.  Controlled Comparative Study
For the controlled study, we use a smaller, fixed architecture of 9 layers with $d_{\text{model}}=1024$, $n_{\text{head}}=32$, and a vocabulary size of 65,536, resulting in models of approximately 247.6M parameters. We compare four models:
  • Constructive Frozen (’Model UNICODE 1_9’): Our proposed method. It is grown incrementally in three stages (layers 1-3, then 4-6, then 7-9). The visual Unicode embedding layer is frozen from the start, as are previously trained layers at each new stage. The final model has 247.6M total parameters, of which 209.9M are frozen.
  • Monolithic Trainable (’Model unfrozen (baseline)’): The primary baseline. A standard 9-layer Transformer with a randomly initialized, fully trainable embedding layer. It is trained end-to-end in a traditional, monolithic fashion. All 247.6M parameters are trainable.
  • Monolithic Frozen (’Model frozen UNICODE (baseline)’): The secondary baseline. A standard 9-layer Transformer with fully frozen and Unicode-predefined embedding layer. It is trained end-to-end in a traditional, monolithic fashion. 67.1M (of 247.6M) parameters are frozen.
  • Constructive Trainable (Ablation 1, ’Model unfrozen 1_9’): This model isolates the effect of the growth strategy itself. It is grown identically to our proposed method (in three stages), but starts with a standard, trainable embedding layer. This tests whether constructive growth is effective without the frozen substrate.
  • Constructive Frozen (16-bit Binary, Ablation 2, ’Model 16_bit 1_9’): This model tests the limits of the emergent semantics principle. It uses a minimalist, 16-dimensional binary vector for each token (derived from the token’s integer ID) as its frozen embedding. This vector is projected to $d_{\text{model}}=1024$ via simple repetition. The model is then grown constructively. The final model has 181.6M total parameters, of which 143.9M are frozen.
  • Monolithic Frozen (’Model frozen 16-bit (baseline)’): The third baseline. A standard 9-layer Transformer with fully frozen 16-dimensional binary embedding layer. It is trained end-to-end in a traditional, monolithic fashion. 1.0M (of 181.6) parameters are frozen.
4.3.  Training Datasets and Protocol
The training corpus for all experiments consists of a subset of English Wikipedia combined with SFT datasets (SQuADv2, CommonsenseQA, ARC), totaling approximately 4B tokens. Models are trained using the AdamW optimizer. For the constructive models, training proceeds in stages (e.g., layers 1-3 are trained until convergence, then frozen; then layers 4-6 are added and trained, etc.). For the monolithic baseline, all 9 layers are trained simultaneously from the start.
@@ 5.1.  Results: Progressive Layer-Wise Growth
Figure 4: t-SNE visualization of frozen input token embeddings for ’abs-bvv-6’. In the this frozen embeddings, semantic groups (e.g., numbers, professions, animals) located randomly and without clustering. This fully corresponds to the initial condition of an absence of semantics in the visually precomputed embedding layer.
Figure 5: t-SNE visualization from layer 0 to layer 5 of attention projections (v_proj and o_proj), and transformer block outputs for ’abs-bvv-6’ with frozen input embedding layer. In contrast to the input layer, clear semantic clusters emerge and sharpen with depth, demonstrating that meaning is constructed compositionally by the network.
5.2.  Results: Controlled Comparative Study
The controlled comparative study was designed to rigorously evaluate our constructive learning method against traditional monolithic training and key ablations. The results reveal a clear trade-off between optimization on the training objective (loss) and performance on downstream reasoning tasks (MMLU).
As shown in Figure 6, all models exhibit stable convergence. The monolithic baseline, with all its parameters being trainable, unsurprisingly achieves the lowest final training loss. In contrast, the constructive models, particularly those with frozen components, plateau at a slightly higher loss. This is an expected outcome for two key reasons: first, a significant portion of their parameters are not directly optimized for the training objective; second, and critically, this controlled study intentionally omits the LoRA-based holistic tuning stages used in our large-scale experiment. This was done to strictly isolate the raw effect of the layer-stacking methodology.
The key finding of this study is presented in Figure 7. Our proposed ’Model UNICODE 1_9’ outperforms the monolithically trained baseline on the MMLU benchmark, achieving a final score approximately 5 percentage points higher. This result supports hypothesis that decoupling structural and semantic representation allows the Transformer’s deeper layers to specialize more effectively in high-level reasoning.
Figure 6: Training dynamics during progressive layer-wise growth for ’Model unfrozen (baseline)’, ’Model frozen UNICODE (baseline)’, ’Model frozen 16_bit (baseline)’, ’Model UNICODE 1_9’, ’Model unfrozen 1_9’ and ’Model 16_bit 1_9’. Each loss spike marks the stacking of a new layer group, followed by rapid convergence.
Figure 7: MMLU and ARC-E metric dynamics during progressive layer-wise growth for ’Model unfrozen (baseline)’, ’Model frozen UNICODE (baseline)’, ’Model frozen 16_bit (baseline)’, ’Model UNICODE 1_9’, ’Model unfrozen 1_9’ and ’Model 16_bit 1_9’.
Figure 8: C-SENSE metric dynamics during progressive layer-wise growth for ’Model unfrozen (baseline)’, ’Model frozen UNICODE (baseline)’, ’Model frozen 16_bit (baseline)’, ’Model UNICODE 1_9’, ’Model unfrozen 1_9’ and ’Model 16_bit 1_9’.
Figure 9: t-SNE visualization for the first and the last layer attention projections (v_proj and o_proj), and transformer block outputs for ’Model unfrozen (baseline)’ with trainable embedding layer. In this picture semantic groups (e.g., numbers, professions, animals) tend to make clusters.
Figure 10: t-SNE visualization for the first and the last layer attention projections (v_proj and o_proj), and transformer block outputs for ’Model UNICODE 1_9’ with frozen embedding layer. In this picture semantic groups (e.g., numbers, professions, animals) tend to make clusters after frozen embedding layer demonstrating emergence of semantic structure.
Figure 11: t-SNE visualization of frozen input token embeddings for ’Model unfrozen (baseline)’ (clustering, left) and ’Model UNICODE 1_9’ (no semantic clusters by design, right).
The ablation studies provide further crucial insights. The ’Model unfrozen 1_9’, which is grown constructively but uses standard trainable embeddings, fails to match the performance of either the monolithic baseline or our proposed model on MMLU. This indicates that the constructive growth strategy alone is insufficient; the stability of the frozen substrate is a critical component for achieving superior reasoning. Remarkably, the ’Model 16_bit 1_9’, despite using extremely impoverished 16-bit binary embeddings, still achieves competitive performance, underscoring the Transformer’s powerful ability to induce semantic structure from even minimal, consistent input signals.
These performance differences are mirrored in the models’ internal representations. Figure 11 confirms the initial state: the baseline’s input embeddings are semantically clustered after training, while our model’s are not semantically clustered by design. However, visualizations of the final layers (Figure )10) show that rich semantic clusters (e.g., numbers, professions) clearly emerge within the deeper layers of ’Model UNICODE 1_9’, rivaling the structure seen in the baseline (Figure 9). This provides direct visual evidence that semantics are being constructed compositionally within the network, rather than being learned at the input layer.
In summary, this controlled study demonstrates that constructive growth on a frozen substrate is not only a viable training strategy but can lead to superior reasoning performance compared to traditional monolithic training, by allowing the model to more effectively specialize its deeper layers for semantic composition.
@@ 6.1.  The Frozen Embedding as a Universal Docking Port
The key enabler for our methods is the shared, frozen representational substrate. It acts as a universal standard, an "API" or "docking port" for neural components. When all models speak the same fundamental input/output language of visual forms, their higher-level learned knowledge (the transformations within their layers) becomes interoperable.
@@ 6.3.  Implications for a Modular AI Ecosystem
  • Resource Efficiency and Specialization: Organizations could train smaller, expert models on proprietary or specialized data.
@@
  • Synergy with Model Quantization: Our approach, particularly with the minimalist 16-dimensional binary embeddings, introduces a "quantization-native" input layer. Standard models learn continuous, high-precision embeddings that must be aggressively quantized post-training. In contrast, our method is built upon an inherently discrete foundation: the embedding vectors themselves are derived from binary, monochrome glyph images. This creates a natural synergy with quantization techniques, as the model is trained from the outset to compose meaning from low-precision, discrete inputs, rather than learning continuous representations that must be aggressively quantized post-hoc. This invites future research into whether a discrete representational substrate can facilitate more efficient and performant quantization of the entire model.
@@ 7.  CONCLUSION
Building on the foundation of frozen visual embeddings, we have demonstrated novel and efficient paradigm for progressive layer-wise growth. This work reframes the challenge of scaling AI from a monolithic endeavor to a constructive and modular one. By establishing a fixed, universal representational substrate, we unlock a new design space for creating powerful, flexible, and efficient AI systems. This "constructive learning" paradigm offers a promising path toward a more sustainable and collaborative future for artificial intelligence.